A initial velocity 6.0 m/s East accelerates at 4.0 m/s^2 east. How long will it take the motorcycle to reach the final velocity
1
Expert's answer
2020-11-30T14:55:09-0500

Let the final velocity be "v_f". Then, by definition, the acceleration is:


"a = \dfrac{v_f-v_i}{t}"

where "v_i = 6m\/s" and "t" is a time taken to reach the final velocity. Thus, obtain:


"t = \dfrac{v_f-v_i}{a}\\
t = \dfrac{v_f-6}{4}\space [s]"

Substituting any "v_f" (in m/s) into this formula, obtain the time in seconds.


Answer. "t = \dfrac{v_f-6}{4}\space [s]."